OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where A Subcript 1 = 15 and A subcript15 = –41

OpenStudy (anonymous):

a1 = 15 = a + (n - 1)d SO, a + (1 -1)d = 15 a = 15 Similarly, a15 = -41 = a + (15 - 1)d a + 14d = -41 14d = -41 - a 14d = -41 - 15 14d = -56 d = -4 So here we are to find 41st term ie a41, so here n = 41 a41 = a + (41 - 1)d = 15 + 40(-4) a41 = 15 - 160 = -145 is the 41st term...